OpenStudy (anonymous):

Part 1: Solve the inequality -5x + 8 <= x - 4 and show all of your work. You can use <= to represent the less than or equal to symbol. Part 2: Use complete sentences to describe the graph of the solution from Part 1. Be sure to include the type of circle and which direction the shading will be. Answer

OpenStudy (karatechopper):

part 1

OpenStudy (anonymous):

ok so, you have to collect all X terms on one sde and all constant terms on the other side.

OpenStudy (karatechopper):

-5x+8<=x-4 -5x+8-8<=x-4-8 -5x<=x-12 -5x-x<=x-x-12 -6x<=-12 -6x/-6=>-12/-6 x=>2

OpenStudy (karatechopper):

no u just write in words about what u would do step by step on what ur graph would look like and how u would graph it

OpenStudy (karatechopper):

first word was now

OpenStudy (anonymous):

kould u do that too>

OpenStudy (karatechopper):

well u have to do ur own work no way i am doin ur homework for when i am studying for my finals

OpenStudy (anonymous):

kk thanks anyway ugh.. ok